USA Canada
oranges salmonberries oranges salmonberries
12 0 6 0
3 3 3 3
0 4 0 6


a. Which country has the comparative advantage in producing salmonberries? Which one has the comparative advantage in producing oranges? Explain.

b. Suppose that each country is currently producing 3 thousand bushels of salmonberries and 3 thousand bushels of oranges. Show that both can be better off if they specialize in producing one good and then engage in trade.
1
Expert's answer
2019-05-31T11:51:18-0400

a) Canada has the comprative advantage in salmonberries.

Alternative cost of thousand bushels of salmonberry there is 1 thousand bushels of orange (the country produces either 6 thousand bushels of oranges or 6 thousand bushels of salmonberries), while in the USA 1 thousand bushels of salmonberry=3 thousand bushels of oranges (the country produces either 12 thousand bushels of oranges or 4 thousand bushels of salmonberries).

The USA has the comprative advantage in oranges, because its alternative cost of thousand bushels of orange is 1/3 thousand bushels of salmonberry.

b) Consider the case when both countries specialize in producing one good according to their comparative advantages and then engage in trade. If Canada produces totally 6 thousand bushels of salmonberries, it can sell 3 thousand bushels of salmonberries in the USA and get 3*3=9 thousand bushels of oranges (instead of 3 thousand bushels of oranges). Totally Canada will gain 9-3=6 thousand bushels of oranges.

Similarly, if the USA produces 12 thousand bushels of oranges, it can sell 9 thousand bushels of oranges in Canada and get 9 thousand bushels of salmonberries. Totally the USA will gain 9-3=6 thousand bushels of salmonberries.